OpenStudy (help!!!!):

6. A car dealer sells a SUV for $39,000, which represents a 25% profit over the cost. What was the cost of the SUV to the dealer? A. $29,250 B. $31,200 C. $32,500 D. $33,800 E. $33,999

OpenStudy (mathstudent55):

The car cost the dealer x. He sells it for a 25% profit. That means he sells it for x + 25%x. But x + 25%x = x + 0.25x = 1.25x So, 1.25x = 39,000 Now solve for x, his cost.